OpenStudy (anonymous):

Anastasia uses the equation p = 0.7(rh + b) to estimate the amount of take-home pay, p, for h hours worked at a rate of r dollars per hour and any bonus received, b. What is an equivalent equation solved for h? h = (– b)÷ r h = – b ÷ r h = ÷ r – b h = ÷ r

OpenStudy (welshfella):

p = 0.7(rh + b) divide both sides by 0.7 rh + b = p/0.7 subtract b from both sides rh = (p/0.7) - b can you finish this off?

OpenStudy (anonymous):

would it be c

OpenStudy (anonymous):

im lost now

OpenStudy (welshfella):

none of the options is correct

OpenStudy (welshfella):

you need to divide both sides by r

OpenStudy (welshfella):

h = (p/0.7r) - b/r or h = p - 0.7b -------- 0.7r